OpenStudy (anonymous):

y varies jointly with x and square of n. When y=25, x=3 and x=4. Find the value of y when x=5 and n=10

OpenStudy (gorv):

y=a*x*n^2 a= constant 25=a*3*4^2 25=a*3*16 25=a*48 find a

OpenStudy (gorv):

put in y=a*x*n^2